OpenStudy (anonymous):

How would the equation r = 2 - 1.5sin(delta) looks like? ( Polar equations)

OpenStudy (anonymous):

Try converting into cartesian, by \[r = \sqrt{x ^{2}+y ^{2}}\] and \[\sin \theta =\frac{ y }{r}\] May be of help..

OpenStudy (loser66):

take a look at http://www.wolframalpha.com/input/?i=r+%3D+2-1.5sin%28theta%29

OpenStudy (anonymous):

The graph of loser66 is good..

OpenStudy (anonymous):

I got it. Thank you :)